Jill is 5 years younger than his brother. Ten years from now, the sum of their ages will be 45. Find Jills current age.
Answer by Cromlix(4381)

Hi there,
Jill is 5 years younger than his brother.
J = B - 5
Ten years from now, the sum of their ages will be 45.
(J + 10) + (B + 10) = 45
J + 10 + B + 10 = 45
Replace J with B - 5
B - 5 + 10 + B + 10 = 45
Collect like terms:
2B = 45 + 5 - 20
2B = 30
B = 15
Jill's current age = 10 years old
Brother's current age = 15 years old.
Hope this helps :-)